Suppose that an annuity will provide for 20 annual payments of 1240 dollars, with the first payment coming 11 years from now. If the nominal rate of interest is 10 percent convertible monthly, what is the present value of the annuity?
Answer by Theo(13342)

the nominal interest ratee is 10% per year.
the effectives interest rate is (1 + .10/12) ^ 12 = (1.1047130674 - 1) * 100 = 10.47130674% per year.
the present value of the annuity 10 years from now is 10225.92263.
the present value of that now is 3777.527116.

here's what it looks like in excel.




eoy means end of year.

let me know if you have any questions.
theo